# RLCAD: Reinforcement Learning Training Gym for Revolution Involved CAD Command Sequence Generation

### v1

@@ 6.2 B-Rep Modeling Comparison
Figure 5: Comparison of different GNN architectures on reconstruction tasks. GTN outperforms GCN and GAT in reconstructing fine geometric details.
We compare three graph neural network (GNN) architectures for reconstructing B-Rep models:
@@
    Graph Convolution Network (GCN): Uses fixed-weight neighborhood aggregation, which is computationally efficient but struggles with capturing heterogeneous geometric details.
@@
    Graph Attention Network (GAT): Incorporates learnable attention mechanisms; however, its single-head attention limits multi-scale feature integration.
@@
    Graph Transformer Network (GTN): Utilizes multi-head self-attention with edge-aware positional encoding, effectively capturing long-range dependencies and global topological constraints.
Table 1 and Figure 5 illustrate that GTN outperforms both GCN and GAT. In particular, it achieves the highest IoU (0.7326), a 0.5% improvement over GCN and a 3.1% gain over GAT. The performance improvement is attributed to the effective processing of relational information between nodes in the graph and the successful capture of long-range dependencies.
Table 1: Performance comparison of GNN architectures on reconstruction tasks. Higher values for IoU, COV, and NC indicate better performance, while lower values for MMD-CD and JSD indicate better.
@@
6.3 Training Efficiency Comparison
We compared our gym environment with Fusion 360 Gym [11] over 10,000 training steps. Using parallel computation (1, 8, and 16 environments), we evaluated both memory consumption and execution time per step. Tables 2 shows that Fusion 360 Gym consumes substantially more memory. The frequent use of the revert operation in Fusion 360 introduces a risk of memory overflow. In contrast, our gym employs a stable mark-and-revert mechanism that maintains steady memory usage without significant fluctuations. In addition, by replacing Fusion 360’s network communication mechanism with multi-process parallelism and shared memory architecture, we achieve significant reduction in data transmission latency between reinforcement learning algorithm and the gym. We get up to $39X$ speedup compared with Fusion 360 Gym, as shown in Table 3.
Table 2: Comparison of memory usage between Fusion 360 and our gym (unit: MB).
Fusion 1-env 8-env 16-env
Initial Memory 1203.6 9.3 10.4 9.5
Final Memory 6362.7 10.2 11.8 10.3
Table 3: Comparison of execution time per step (unit: ms) and achieved speedup.
Fusion 1-env 8-env 16-env
Execution Time 1674 198 52.5 42.18
Speedup - 8.45 31.89 39.69
@@ 6.4 Training Strategy
Our training strategy consists of two phases: feature extraction and command prediction. Since the command encoding retains only face IDs (without explicit geometric details), the pre-trained command generation network loses effectiveness on new geometries. Therefore, we reinitialize the command generation network for each new input. For the feature extraction network, we employ curriculum learning to enhance generalization. We construct a training dataset of 500 geometries, progressively train on the dataset ordered by increasing complexity and update weights from simple models as initialization to complex ones. A comparison with a case-by-case training strategy (where each geometry is trained independently) confirms that curriculum learning leads to superior overall performance, as summarized in Table 4.
@@
Case by case 0.8354 0.8544 0.0165 0.1307 0.8638
Curriculum 0.8757 0.8692 0.0139 0.1111 0.8812
@@ 6.5 Reward Ablation Study
We perform an ablation study on the reward function, which guides the RL agent via cumulative feedback. Table 5 demonstrates that combining IoU, MMD, and NC yields the best performance across all metrics.
Table 5: Reward function ablation study. The combination of IoU, MMD, and NC produces superior results.
@@
6.6 Comparison with Imitation Learning
We further compare our method against an imitation learning approach from Fusion 360 Gallery [11]. Fusion 360 Gallery extracts continuous geometric parameters from B-Rep models and uses imitation learning to generate command sequences, followed by a confidence-based local search for optimization. Fusion 360 Gallery originally limited to extrusion operations, we extend it to include revolve operations. For comparison with Fusion 360 Gallery, we select 9K CAD pairs of B-Rep and command sequence from ABC dataset [27] as the training set, while our curriculum learning algorithm requires the training dataset of 500 B-Rep. The test set of 500 B-Rep are selected from ABC dataset for both our method and Fusion 360 Gallery. Our method achieves higher scores on all evaluation metrics (see Table 6) and produces visually more accurate reconstructions, as illustrated in Figure 6.
Figure 6: Qualitative comparison with Fusion 360 Gallery. Our method can generate more precise command sequence from B-Rep.
Table 6: Quantitative results on B-Rep-based reconstruction tasks. Our method significantly outperforms Fusion 360 Gallery across all metrics.
@@ 6.6 Comparison with Imitation Learning
Fusion 360 Gallery 0.7393 0.6502 0.0678 0.3030 0.7640
Ours 0.8757 0.8692 0.0139 0.1111 0.8812
@@ 6.7 Failure Case
We categorize the cases that our method currently cannot handle into three types. The first type, as shown in Figure 7a, consists of complex models with smart features. Our method fails to learn the slight feature in hidden space. The second type arises from the limitations of the supported operations of our training gym. For example, models with fillet and spline surface, as shown in Figure 7b. We will support such operations in the future. The third type is related to the reward design. For example, in Figure 7c, since the sphere is the primary geometric feature and the reward score exceeds 0.99, subsequent details such as extrude receive minor reward variations, which has minimal impact on the network training.
Figure 7: Examples of failure cases categorized into three types: (a) complex models with numerous slight features; (b) models requiring unsupported operations (e.g., fillet and spline surface); (c) reward mechanism failure.
@@ 7.1 Conclusion
Most current CAD model generation algorithms are based on supervised learning methods, where the validity of the generated geometries cannot be effectively verified and fed back to the training network. We are the first to present a geometric engine-based CAD model training gym and introduce an online RL algorithm based on the gym to CAD model generation. Our RL algorithm has added support for revolution operation, enhancing the capability of generating comand sequence for complex models. Experiments demonstrate our method can significantly improve the precision of command sequence generation compared with supervised learning methods.
@@ 7.2 Future Work
Through extensive experiments, it has been found that our method still has shortcomings in detailed feature learning and reward design. Additionally, the current training gym only supports extrude and revolve operations. In the next phase, we will focus on developing a more detail-friendly feature extraction network, designing reward functions that better balance primary and detailed features, and expanding the gym to support more commonly used modeling operations.

### v2

@@ 6.3 B-Rep Modeling Comparison
Figure 5: Comparison of different Graph Neural Network (GNN) architectures and UV-Net on reconstruction tasks. It indicates that the UV-Net outperforms the Graph Convolutional Network (GCN), Graph Attention Network (GAT), and Graph Transformer Network (GTN) in reconstructing fine geometric details.
We compare three graph neural network (GNN) architectures and UV-Net for reconstructing B-Rep models:
@@
    Graph Convolution Network (GCN): Use fixed-weight neighborhood aggregation, which is computationally efficient but struggles with capturing heterogeneous geometric details.
@@
    Graph Attention Network (GAT): Incorporate learnable attention mechanisms; however, its single-head attention limits multi-scale feature integration.
@@
    Graph Transformer Network (GTN) [46]: Utilize multi-head self-attention with edge-aware positional encoding, effectively capturing long-range dependencies and global topological constraints.
  • 4.
    UV-Net: Exploit the UV parametric domain of curves and surfaces for geometry modeling and adjacency graphs for explicit topological representation. This approach effectively combines convolutional neural networks and graph neural networks to significantly enhance representation capability.
To enhance the discriminative power of the ablation analysis, we curated $0.5k$ complex models from the test set to construct the ablated dataset. As shown in Table 3, UV-Net outperforms GCN, GAT, and GTN across all five evaluation metrics on this dataset. In particular, it achieves a significant improvement in IoU compared to the previous best-performing method, GTN. The model comparison in Figure 5 further substantiates this finding, demonstrating UV-Net’s superior capability in learning latent feature representations of B-Rep models, which results in a closer fit to the ground truth.
Table 3: Performance comparison of GNN architectures and UV-Net on reconstruction tasks. The direction of the arrows indicates the direction of better performance. It shows that UV-Net outperforms GCN, GAT, and GTN across all five evaluation metrics.
@@
UV-Net 0.8207 0.7069 0.0104 0.2303 0.7839
@@ 6.4 Training strategy
Our training strategy consists of two phases: feature extraction and command prediction. Since the command encoding retains only face IDs (without explicit geometric details), the pre-trained command generation network loses effectiveness on new geometries. Therefore, we reinitialize the command generation network for each new input. For the feature extraction network, we perform self-supervised pre-training on $189k$ CAD models and employ curriculum learning to enhance generalization. We first rank the CAD models in order of increasing complexity based on the number of faces, and then train them sequentially from low to high complexity, updating the model weights along the way. A comparison with a case-by-case training strategy (where each geometry is trained independently) confirms that curriculum learning leads to superior overall performance, as summarized in Table 4.
@@
Case by case 0.8142 0.7567 0.0194 0.2437 0.7754
Curriculum 0.8624 0.7725 0.0078 0.2076 0.7923
@@ 6.5 Reward Ablation Study
Table 5: Reward function ablation study. It shows that introducing NR significantly improves evaluation metrics of IoU, COV, JSD and NC, while metric of MMD-CD shows no significant degradation. Therefore, we choose the combination of IoU, MMD, NC, and NR as our final reward configuration.
@@
$\mathrm{IoU+MMD+NC+NR}$ 0.8712 0.7408 0.0106 0.2178 0.7908
Leveraging the ablated dataset curated in Section 6.3, we perform ablation studies to systematically analyze the impact of different reward function combinations on experimental results. Table 5 shows the outcomes of these studies. Reward functions such as IoU, MMD, and NC are used to measure geometric similarity. NR is used to capture the cosine similarity of high-dimensional B-Rep features. The experimental results show that introducing NR significantly improves evaluation metrics of IoU, COV, JSD and NC, meanwhile, metric of MMD-CD shows no significant degradation. Therefore, we choose the combination of IoU, MMD, NC, and NR as our final reward configuration.
6.6 Result Comparison of Reconstruction
Table 6: Quantitative results on the B-Rep-based reconstruction task. Our method significantly outperforms Fusion 360 Gallery across all metrics.
@@ 6.6 Result Comparison of Reconstruction
Fusion 360 Gallery 0.5678 0.4352 0.4769 0.4968 0.6299
Ours 0.9001 0.8206 0.0180 0.1945 0.8227
Table 7: Quantitative comparison with cadrille and CAD-Recode on the reconstruction task. Our method significantly outperforms these two methods across all metrics.
[…]
cadrille 0.7603 0.7260 0.0577 0.4904 0.6001
CAD-Recode 0.9264 0.8736 0.0043 0.0885 0.8205
Ours 0.9779 0.9836 0.0021 0.0618 0.8799
We select three representative learning based CAD model generation algorithms for comparison.
  • 1.
    Fusion 360 Gallery [11] extracts continuous geometric parameters from B-Rep models and uses imitation learning to generate command sequences, followed by a confidence-based local search for optimization.
  • 2.
    CAD-Recode [47] translates a point cloud into Python code by leveraging a relatively small language model as a decoder and combined with a lightweight point cloud projector. The method is trained on a self-constructed, million-scale training dataset composed exclusively of sketch-extrusion operations.
  • 3.
    cadrille [48] proposes a multi-modal CAD reconstruction model that can simultaneously process three input modalities, including point clouds, images, and text. The method adopts a two-stage pipeline: supervised fine-tuning (SFT) on large-scale procedurally generated data, followed by reinforcement learning fine-tuning using online feedback.
Currently, most supervised learning methods are trained on datasets that include only sketch-extrusion operations. To overcome this limitation, we extend the output modules of Fusion 360 Gallery models to support revolve operations. And we retrain the model on our dataset. As illustrated in Figure 6, our approach outperforms Fusion 360 Gallery in completeness and detail preservation. Quantitative results in Table 6 demonstrate substantial improvements across all evaluation metrics.
To validate the generalization capability of our method on unseen datasets, we select two baseline approaches - CAD-Recode (using point cloud as input) and cadrille (using point cloud and image as input) - both pre-trained on the CAD-Recode dataset, and compare their generation results with our method on the CAD-Recode validation dataset. As illustrated in Figure 7 and summarized in Table 7, which present qualitative results and quantitative metrics respectively, our method consistently outperforms both baselines across all evaluation criteria.
Figure 6: Comparison of generation results with Fusion 360 Gallery. It shows our method generates higher-quality results in terms of both completeness and detail.
Figure 7: Comparison of generation results with cadrille and CAD-Recode. It shows our method generates higher-quality results in terms of both completeness and detail.
@@ 6.7 Failure Case
We categorize the cases that our method currently cannot handle into three types. The first type arises from complex model with many-faces. Figure 8a exhibits the generation result of a complex gear model, in which some details is missing. Figure 9 depicts the quantitative relationship between generated model complexity (measured by face count) and various evaluation metrics. For simple models with a low face count, IoU and NC remain relatively high, but gradually decline as complexity increases. COV performs well on models with a low number of faces but declines significantly once the face count exceeds approximately 35–40. JSD, which is initially low, rises rapidly as the face count increases, especially beyond 40 faces. Overall, while the global reconstruction quality of our model remains acceptable beyond 35–40 faces, the fidelity of fine-grained geometric details deteriorates noticeably.
Figure 8: Examples of failure cases categorized into three types: (a) complex model with many faces; (b) models requiring unsupported operations (e.g., fillet and spline surface); (c) special model trimming.
Figure 9: The quantitative relationship between model complexity and evaluation metrics. While the global reconstruction quality of our model remains acceptable beyond 35–40 faces, the fidelity of fine-grained geometric details deteriorates noticeably.
The second type arises from the limitations of the supported operations of our training gym. For example, models with fillet and spline surface, as shown in Figure 8b. We will support such operations in the future.
The third type of failure case arises from the inconsistency between the selected trimmed body and the requirements during model trimming. As shown in Figure 8c, the reconstruction of the ground truth model is completed by clipping a sphere twice: first, a complete sphere is constructed, then a groove based on the extrusion is built at the waist, and a cylinder based on rotation is constructed at the center. Finally, these two entities are cut from the sphere through Boolean operations. But in reality, the outer diameter of the circular ring at the bottom of the waist groove and the height of the rotating surface of the central cylinder are both smaller than the diameter of the sphere. The entities generated by extrusion and rotation are both inside the sphere, so the surface of the sphere remains complete after clipping.
@@ 7.1 Conclusion
Most current CAD model generation algorithms are based on supervised learning methods, where the validity of the generated geometries cannot be effectively verified and fed back to the training network. We are the first to present a geometric engine-based CAD model training gym and introduce an online RL algorithm based on the gym to CAD model generation. Our RL algorithm has added support for revolution operation, enhancing the capability of generating command sequences for complex models. Experiments demonstrate that our method can significantly improve the precision of the command sequence generation compared with supervised learning methods.
@@ 7.2 Future Work
Through extensive experiments, it has been found that our method still has shortcomings in detailed feature learning of complex models and supported operation amount. In the next phase, We will focus on developing detail-preserving feature extraction networks that support a broader range of input modalities, and expanding the gym to support more commonly used modeling operations.